OpenStudy (anonymous):

don't know how to do this please help A bacteria culture begins with 13 bacteria which triple in amount at the end of every hour. How many bacteria exist in the culture at the end of the 10th hour? 13,312 26,624 767,637 2,302,911

OpenStudy (shubhamsrg):

initially,no. is 13 then it becomes 13*3 then (13*3)*3 . . . 10 times so you have a GP like this: 13, (13*3) , (13*3)*3 ..... here a=13 r=3 you need to find sum of this Gp for 10 terms sum = a(r^n -1)/(r-1) just make these substitutions hope this helps.

OpenStudy (anonymous):

= 13 * (3) ^10 = ...

OpenStudy (anonymous):

@kerber123 can you finish it up?

OpenStudy (anonymous):

767,637?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

again ??

OpenStudy (shubhamsrg):

ohh,,am sorry,,i misread the ques..really sorry..

OpenStudy (anonymous):

@shubhamsrg It just proves that your thought so complicated than mine :)

OpenStudy (shubhamsrg):

i'll take that as a compliment!! lol..

OpenStudy (anonymous):

I honestly mean it ;)